OpenStudy (anonymous):

I'll medal and fan pls help Determine whether each equation is true or false. Select True or False for each equation. −5/−1=5 True False ( 0/−4)=0 True False −3/1=−3 True False −8/0=0 True False

iYuko (iyuko):

1: true neg div neg =pos

iYuko (iyuko):

2: true

iYuko (iyuko):

3: true

OpenStudy (anonymous):

Can you explain how it's true/false?

iYuko (iyuko):

2: pos div ned=pos

iYuko (iyuko):

3: neg div pos =neg

OpenStudy (anonymous):

Ok thank you

iYuko (iyuko):

4: is a neg

iYuko (iyuko):

idk bc you cant divide by 0

OpenStudy (anonymous):

false for the 4th one?

iYuko (iyuko):

yes

OpenStudy (anonymous):

Alright, uh does everyone agree? Should I put that?

iYuko (iyuko):

Im not 100% sure on 4th one because you can't divide by 0

OpenStudy (anonymous):

ok that's fine

OpenStudy (anonymous):

\[\frac{ -8 }{ 0 }=undefined\]

OpenStudy (usukidoll):

Since you can't divide by 0, the fraction becomes undefined. THerefore, the statement is false automatically because of that .

iYuko (iyuko):

Thank you @UsukiDoll

OpenStudy (anonymous):

yea :)

OpenStudy (anonymous):

Oh I see thank you

OpenStudy (anonymous):

I guess i'll just medal I yuko idk

iYuko (iyuko):

Medal who YOU think helped you.

OpenStudy (anonymous):

Thanks everyone :)

OpenStudy (usukidoll):

It's like... For example 8/0 we can't divide by 0, so the fraction is undefined. but.. we can have a 0 on the numerator Ex 0/8 = 0 because 0 divided by 8... we're dividing nothing by 8 so the result is 0. For the first and third questions, the explanation is decent because on the first a negative divided by a negative will result in a positive due to the signs canceling out. for the third and negative divided by a positive will still be negative

OpenStudy (anonymous):

Thank you UsukiDoll I understand it much better

OpenStudy (anonymous):

Always remember\[(+)\times(+)=+\]\[(-)\times(-)=+\]\[(-)\times(+)=-\]\[(+)\times(-)=-\]

OpenStudy (anonymous):

Oh ok thanks!

OpenStudy (usukidoll):

ah yes the traditional charts for signs @MARC_ :)

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

np @LoveIt

OpenStudy (johan14th):

Simple rule is if there is a number on top and a zero on the bottom it is undefined (number/ zero) n/0 i use this to remind me no because n/0 is no

OpenStudy (johan14th):

If its zero 0/number it is okay 0/n can be seen as on or okay which would equal to 0

OpenStudy (johan14th):

Sorry if that was confusing